# Citation Sentiment Reflects Multiscale Sociocultural Norms

### v1

@@ Abstract
Modern science is formally structured around scholarly publication, where scientific knowledge is canonized through citation. Precisely how citations are given and accrued can provide information about the value of discovery, the history of scientific ideas, the structure of fields, and the space or scope of inquiry. Yet parsing this information has been challenging because citations are not simply present or absent; rather, they differ in purpose, function, and sentiment. In this paper, we investigate how critical and favorable sentiments are distributed across citations, and demonstrate that citation sentiment tracks sociocultural norms across scales of collaboration, discipline, and country. At the smallest scale of individuals, we find that researchers cite scholars they have collaborated with more favorably (and less critically) than scholars they have not collaborated with. Outside collaborative relationships, higher h-index scholars cite lower h-index scholars more critically. At the mesoscale of disciplines, we find that wetlab disciplines tend to be less critical than drylab disciplines, and disciplines that engage in more synthesis through publishing more review articles tend to be less critical. At the largest scale of countries, we find that greater individualism (and lesser acceptance of the unequal distribution of power) is associated with more critical sentiment. Collectively, our results demonstrate how sociocultural factors can explain variations in sentiment in scientific communication. As such, our study contributes to the broader understanding of how human factors influence the practice of science, and underscore the importance of considering the larger sociocultural contexts in which science progresses.
@@ I Introduction
In the modern era, science has often been assumed to be an objective process [1]. By this descriptor, scholars have variously meant faithfulness to facts, absence of normative commitments, and freedom from personal biases [2]. Yet with an increasing understanding of human psychology has come an acknowledgement that even the most scrupulous of researchers cannot achieve various types of objectivity; for example, (1) the mechanical objectivity of suppressing the universal human propensity to judge and aestheticize, or (2) the aperspectival objectivity of eliminating individual idiosyncracies [3]. The growing appreciation of human subjectivity has motivated a turn from the personal to the community, where objectivity is considered a quality that characterizes a collection or population of studies [4], and hence is a feature of scientific communities and their practices [5]. In communities, objectivity can further be thought of as occurring in degrees, such that any method of inquiry is only objective to the degree that it permits transformative criticism through shared standards, equality of intellectual authority, avenues for criticism, and uptake of criticism [6].
@@
Citation networks have played a key role in the sociology of science and its critical evaluation [27, 24, 28, 29]. Yet, due to various challenges of data accessibility and limitations in algorithmic efficacy, a major gap in knowledge exists: precisely how, or in what way, do we cite? Citations can be differently valued [30]; can be epistemic or procedural [18]; can be positive or negative, indicate either support or contrast, attribute creativity or indicate status as a “classic” [31, 32]; can either represent or misrepresent the work cited [33]; and can be chosen for social reasons (such as name-dropping [33] or soliciting favor from a scholar who might have influence in the review process), for mercantile reasons (such as bartering for a citation in return), and for alignment reasons (such as transmitting a specific self-image, e.g., as part of the mainstream, the avant garde, or a given school) [34, 35]. Because not all citations are created equal, citation analysis could be meaningfully expanded to account for citation type [28, 34]. Although a few early studies attempted such an expansion, they were limited in scope by the lack of computational infrastructure to assess citation type automatically from large databases.
@@
Here we overcome this fundamental limitation in order to examine citation sentiment and its dependence upon the sociocultural factors in which science exists and progresses (Fig. 1). To ensure that our assessment is computationally tractable while remaining sensitive to local cultural norms, we focus on a single discipline—neuroscience—and build a large database of papers from 56,000 last authors published in 185 journals (impact factor $\geq 3$ in 2022). As a candidate field, neuroscience has particular affordances for our purpose: its relative youth allows for the emergence and growth of not-yet-crystallized subdisciplines, a fundamental interdisciplinarity as it draws on earlier-defined disciplines, and a marked intersection of opinions, disciplinary methods, and approaches to study [36, 37]. All of these factors can support a diversity, heterogeneity, and complexity of the space of ideas, and the network of citations among papers espousing those ideas. We chose to focus on citation sentiment (neutral, favorable, or critical), as a kind or type, for functional and pragmatic purposes. Functionally, sentiment can track coalescence (or fragmentation) of people and ideas; pragmatically, sentiment can be measured quantitatively in extensive databases using large language models. Using this approach, we show how citation sentiment varies across people groups, scientific disciplines, and whole countries, providing a paper-trail lens into the socio-cognitive processes of science.
@@
Figure 2: Citation sentiment depends on distance and years since/to first collaboration. (A) Citation sentiment relative to the null model as a function of the collaboration distance between the last authors of the citing and cited works. Shaded regions denote one standard deviation from the mean at a given collaboration distance. (B) Citation sentiment as a function of the number of years since the first collaboration between the two last authors. A negative year implies that at the time of a citation, the two scholars had yet to collaborate but would do so in the future. Shaded regions denote one standard deviation.
@@ II Results
In a large connected citation network of $108909$ papers from the PMC Open Access Subset, we assess citation sentiment as neutral, favorable, or critical (see Supplement). Neutral sentiment is observed in factual phrases or statements, devoid of opinion (e.g., “was reported”, “has been shown”, “can alter”, “can play a role”, “is”, “responds to”). Favorable sentiment is observed in phrases or statements that indicate a positive assessment or evaluation (e.g., adjectival phrases like ”elegantly shown”, “optimized method”, “powerful tools”) or that indicate consensus or consistency (e.g., “converging results”, “conclusion is consistent”). Critical sentiment is observed in phrases or statements that indicate debate (“the focus of debate”), disagreement (“while other studies do X, we argue Y”), difference (“in contrast to a previous report”), distinction (“is not applicable here”), distance (“despite the fact that many studies have evaluated X, very few have evaluated Y”), or limitation (“however, X method limits desirable outcome Y”). In this formulation, then, sentiment captures both social and epistemic alignment by measuring evaluation and intellectual commonality, respectively. After extracting $627108$ citation sentences and parsing them, we observed that $37.3\%$ were favorable, $5.3\%$ were critical, and $57.4\%$ were neutral. Our first question was how these sentiments were aligned with ingroup/outgroup relations and structures of dominance. Specifically, we investigated how sentiment is assigned to the ingroup of collaborators (vs. the outgroup of scholars one has never co-authored with) and to the dominant group of scholars with high h-indices (vs. the less dominant group of scholars with lower h-indices).
@@ II.1 Citation Sentiment is Most Favorable to Collaborators
To examine how sentiment might reflect ingroup/outgroup relations, we measured the collaboration distance: the geodesic distance in the collaboration network between the citer and the citee [38] at the year of citation. We observed that people are most favorable to themselves and collaborators, and most critical to scholars with whom they have not collaborated (Fig. 2A). In subsequent analyses, we removed self-citations to focus on community practices. Note that collaboration distance likely tracks both social distance and intellectual distance. What happens when intellectual distance is held fixed, and social distance changes? To take an initial step toward addressing this question, we isolate citer-citee pairs who would eventually become collaborators. Just prior to collaborating, we find that sentiment is overwhelmingly critical, whereas just after collaborating, critical sentiment plummets and favorable sentiment rises (Fig. 2B). This pattern suggests that the social act of collaboration may drive a transformation from critical to favorable sentiment among scholars. We also notice that citer-citee pairs that will collaborate within the next four years tend to display significantly greater critical sentiment ($46.6\%$ more than expected) than citer-citee pairs that will not collaborate ($1.2\%$ more than expected; two-sided Welch’s $t$-test: $t=91.3$, $p<0.0001$, $df=1003.8$). The presence of this dramatic increase could indicate that critical language in citation may be an early predictor of future collaboration.
Figure 4: Critical citation sentiment varies across disciplines. (A) Mean critical sentiment in a discipline toward collaborators versus non-collaborators. (B) Mean critical sentiment as a function of the benchwork score of the citing paper’s discipline. We define the benchwork score as the estimated proportion of wet lab papers (instead of dry lab papers) in a discipline. Orange corresponds to the sentiment toward non-collaborators, cyan to the sentiment toward collaborators, and grey to the sentiment toward both. Lines are fit using weighted least squares regression; shaded regions represent $95\%$ confidence intervals. (C) Mean critical sentiment as a function of the synthesis score of a citing paper’s discipline. We define synthesis as the proportion of review papers in a discipline. (D) Relationship between the mean critical sentiment and bias. We define bias as the difference between sentiments toward non-collaborators and collaborators. (E) Bias in critical sentiment as a function of a discipline’s benchwork score. (F) Bias in critical sentiment as a function of disciplinary synthesis.
@@ II.3 Citation Sentiment Tracks Disciplinary Differences & Practices
Figure 5: Favorable citation sentiment varies across disciplines. (A) Mean favorable sentiment in a discipline toward collaborators versus non-collaborators. (B) Mean favorable sentiment as a function of the benchwork score of the citing paper’s discipline. We define benchwork score as the estimated proportion of wet lab papers (as opposed to dry lab papers) in a discipline. Orange corresponds to the sentiment toward non-collaborators, cyan to the sentiment toward collaborators, and grey to the sentiment toward both. Lines are fit using weighted least squares regression; shaded regions represent $95\%$ confidence intervals. (C) Mean favorable sentiment as a function of the synthesis score of a citing paper’s discipline. We define synthesis as the proportion of review papers in a discipline. (D) Relationship between the mean favorable sentiment and bias. We define bias as the difference between sentiments toward non-collaborators and collaborators. (E) Bias in favorable sentiment as a function of a discipline’s benchwork score. (F) Bias in favorable sentiment as a function of disciplinary synthesis.
The previous two sections suggest that citation sentiment tracks social factors at the level of inter-personal differences (h-index) and relationships (collaboration). We expand out from dyads to larger social groups in the form of scientific disciplines, with their own broader social norms, cultures [39, 40], and identities [41]. Moreover, as so-called communication complexes [42] with their own rhetoric [43], writing styles [44], and communicative processes [42], we expect disciplines to differ in citation sentiment and in the prevalence of critical vs. favorable sentiment to collaborators and strangers. Consistent with our expectations, we see that critical sentiment varies by discipline (Fig. 4A), with most disciplines displaying less critical sentiment to collaborators than to non-collaborators (Fig. 4A, the sector above the unity line). Interestingly, we observe that the degree of critical sentiment in a discipline overall is positively correlated with sentiment bias; that is, the more critical the discipline in general, the more likely the discipline will more critically cite non-collaborators than collaborators (Fig. 4D; linear regression adjusted $R^{2}=0.294$; F-test $F(1,25)=11.81$, $p<0.01$). Similarly, we see that favorable sentiment varies by discipline (Fig. 5A), with all disciplines displaying more favorable sentiment to collaborators than to non-collaborators (Fig. 5A, the sector below the unity line). The degree of favorable sentiment in a discipline overall is significantly correlated with sentiment bias (Fig. 5D; adjusted $R^{2}=0.498$, $F(1,25)=26.77$, $p<0.0001$).
@@
We expect disciplines offering lower-level explanations from benchwork to have less critical sentiment because that scale of explanation is more accepted and valued. By contrast, we expect disciplines offering higher-level explanations to more frequently resort to the use of sentiment to signal valuation of the scientific work. Consistent with our expectations, we observe a significant negative trend between critical sentiment and the benchwork score (linear regression adjusted $R^{2}=0.392$; F-test $F(1,25)=17.76$, $p<0.001$; Fig. 4B, grey). Interestingly, the effect holds both when citing non-collaborators (adjusted $R^{2}=0.380$, $F(1,25)=16.91$, $p<0.001$) and when citing collaborators ($R^{2}=0.302$, $F(1,25)=12.22$, $p<0.01$). Similarly, we observe a significant negative trend between favorable sentiment and the benchwork score (linear regression adjusted $R^{2}=0.406$; F-test $F(1,25)=18.79$, $p<0.001$; Fig. 5B, grey). Again the effect holds when citing non-collaborators (adjusted $R^{2}=0.392$, $F(1,25)=17.74$, $p<0.001$), but is significantly mitigated when citing collaborators ($R^{2}=0.002$, $F(1,25)=1.06$, $p=0.31$). Further, we observe a significantly negative trend between the benchwork score and bias in favorable sentiment towards collaborators (adjusted $R^{2}=0.369$, $F(1,25)=16.18$, $p<0.001$; Fig. 5E). This negative trend is not seen for critical sentiment (linear fit and F-test between benchwork score and critical sentiment bias $R^{2}=0.054$, $F(1,25)=2.50$, $p=0.13$; Fig. 4E).
@@
We define synthesis to be the proportion of review papers in a discipline. Consistent with our expectations, we observe a significant negative correlation between critical sentiment and synthesis (linear regression adjusted $R^{2}=0.607$; F-test $F(1,25)=41.21$, $p<0.0001$; Fig. 4C, grey). The effect holds when citing non-collaborators (adjusted $R^{2}=0.602$, $F(1,25)=40.41$, $p<0.0001$), but is mitigated when citing collaborators ($R^{2}=0.277$, $F(1,25)=10.95$, $p<0.01$). Further, we observe that disciplines with greater synthesis display less bias in critical sentiment towards non-collaborators, and disciplines with less synthesis display more bias (linear fit and F-test between synthesis and critical sentiment bias $R^{2}=0.201$, $F(1,25)=7.55$, $p<0.05$; Fig. 4F). Similarly, we observe a significant negative correlation between favorable sentiment and synthesis (linear regression adjusted $R^{2}=0.639$; F-test $F(1,25)=47.06$, $p<0.0001$; Fig. 5C, grey). Again the effect holds when citing non-collaborators (adjusted $R^{2}=0.609$, $F(1,25)=41.53$, $p<0.0001$), and is similarly mitigated when citing collaborators ($R^{2}=0.115$, $F(1,25)=4.37$, $p<0.05$). We observe a significant negative relationship between synthesis and bias in favorable sentiment towards collaborators (adjusted $R^{2}=0.473$, $F(1,25)=24.36$, $p<0.0001$; Fig. 5F). Note that our null model (see Methods) accounts for the fact that review papers in general express less sentiment ($28.56\%$ favorable, $4.69\%$ critical) compared to research papers ($40.52\%$ favorable, $5.52\%$ critical).

### v2

@@ Abstract
Modern science is formally structured around scholarly publication, where scientific knowledge is canonized through citation. Precisely how citations are given and accrued can provide information about the value of discovery, the history of scientific ideas, the structure of fields, and the space or scope of inquiry. Yet parsing this information has been challenging because citations are not simply present or absent; rather, they differ in purpose, function, and sentiment. In this paper, we investigate how critical and favorable sentiments are distributed across citations, and demonstrate that citation sentiment tracks sociocultural norms across scales of collaboration, discipline, and country. At the smallest scale of individuals, we find that researchers cite scholars they have collaborated with more favorably (and less critically) than scholars they have not collaborated with. Outside collaborative relationships, higher h-index scholars cite lower h-index scholars more critically. At the mesoscale of disciplines, we find that wetlab disciplines tend to be less critical than drylab disciplines, and disciplines that engage in more synthesis through publishing more review articles tend to be less critical. At the largest scale of countries, we find that greater individualism (and lesser acceptance of the unequal distribution of power) is associated with more critical sentiment. Collectively, our results demonstrate how sociocultural factors can explain variations in sentiment in scientific communication. As such, our study contributes to a broader understanding of how human factors influence the practice of science, and underscores the importance of considering the larger sociocultural contexts in which science progresses.
@@ I Introduction
In the modern era, science has often been assumed to be an objective process [1]. By this descriptor, scholars have variously meant faithfulness to facts, absence of normative commitments, and freedom from personal biases [2]. Yet with an increasing understanding of human psychology has come an acknowledgment that even the most scrupulous of researchers cannot achieve various types of objectivity; for example, (1) the mechanical objectivity of suppressing the universal human propensity to judge and aestheticize, or (2) the aperspectival objectivity of eliminating individual idiosyncracies [3]. The growing appreciation of human subjectivity has motivated a turn from the personal to the community, where objectivity is considered a quality that characterizes a collection or population of studies [4], and hence is a feature of scientific communities and their practices [5]. In communities, objectivity can further be thought of as occurring in degrees, such that any method of inquiry is only objective to the degree that it permits transformative criticism through shared standards, equality of intellectual authority, avenues for criticism, and uptake of criticism [6].
@@
Citation networks have played a key role in the sociology of science and its critical evaluation [27, 24, 28, 29]. Yet, due to various challenges of data accessibility and limitations in algorithmic efficacy, a major gap in knowledge exists: precisely how, or in what way, do we cite? Citations can be differently valued [30]; can be epistemic or procedural [18]; can be positive or negative, indicate either support or contrast, attribute creativity or indicate status as a “classic” [31, 32]; can either represent or misrepresent the work cited [33]; and can be chosen for social reasons (such as name-dropping [33] or soliciting favor from a scholar who might have influence in the review process), for mercantile reasons (such as bartering for a citation in return), and for alignment reasons (such as transmitting a specific self-image, e.g., as part of the mainstream, the avant garde, or a given school) [34, 35]. Because not all citations are created equal, citation analysis could be meaningfully expanded to account for citation type [28, 34] and context [36]. Although a few early studies attempted such an expansion, most methods are limited in performance [37] and few use deep learning [38, 39] or large language models [40] to assess citation sentiment in a more comprehensive and efficient manner.
@@
Here, we overcome this fundamental limitation by leveraging large language models to classify citation sentiment while examining its dependence upon the sociocultural factors in which science exists and progresses (Fig. 1). To ensure that our assessment is computationally tractable while remaining sensitive to local cultural norms, we focus on a single discipline—neuroscience—and build a large database of papers from $56,034$ last authors published in $181$ journals (impact factor $\geq 3$ in 2022). As a candidate field, neuroscience has particular affordances for our purpose: its relative youth allows for the emergence and growth of not-yet-crystallized subdisciplines, a fundamental interdisciplinarity as it draws on earlier-defined disciplines, and a marked intersection of opinions, disciplinary methods, and approaches to study [41, 42]. All of these factors can support a diversity, heterogeneity, and complexity of the space of ideas, and the network of citations among papers espousing those ideas. We chose to focus on citation sentiment (neutral, favorable, or critical), as a kind or type, for functional and pragmatic purposes. Functionally, sentiment can track coalescence (or fragmentation) of people and ideas; pragmatically, sentiment can be measured quantitatively in extensive databases using large language models. Using this approach, we show how citation sentiment varies across people groups, scientific disciplines, and whole countries, providing a paper-trail lens into the socio-cognitive processes of science.
@@
Figure 2: Citation sentiment depends on distance and years since (or to) first collaboration. (A) Citation sentiment relative to the null model as a function of the collaboration distance between the last authors of the citing and cited works. Shaded regions denote one standard deviation from the mean at a given collaboration distance. (B) Citation sentiment as a function of the number of years since the first collaboration between the two last authors. A negative year implies that at the time of a citation, the two scholars had yet to collaborate but would do so in the future. Shaded regions denote one standard deviation.
@@ II Results
In a large connected citation network of $108909$ papers from the PMC Open Access Subset, we assess citation sentiment as neutral, favorable, or critical (see Supplement). Neutral sentiment is observed in factual phrases or statements, devoid of opinion (e.g., “was reported”, “has been shown”, “can alter”, “can play a role”, “is”, “responds to”). Favorable sentiment is observed in phrases or statements that indicate a positive assessment or evaluation (e.g., adjectival phrases like ”elegantly shown”, “optimized method”, “powerful tools”) or that indicate consensus or consistency (e.g., “converging results”, “conclusion is consistent”). Critical sentiment is observed in phrases or statements that indicate debate (“the focus of debate”), disagreement (“while other studies do X, we argue Y”), difference (“in contrast to a previous report”), distinction (“is not applicable here”), distance (“despite the fact that many studies have evaluated X, very few have evaluated Y”), or limitation (“however, X method limits desirable outcome Y”). In this formulation, then, sentiment captures both social and epistemic alignment by measuring evaluation and intellectual commonality, respectively.
After measuring the sentiment of $627108$ extracted citation sentences, counted by the number of citation locations in a sentence, we assign sentiment values to each of the citer-citee pairs and aggregate them (see Methods and Supplement for further details). We observe that $42.21\%$ are favorable, $7.94\%$ are critical, and $49.85\%$ are neutral. Our first question is how these sentiments align with ingroup/outgroup relations and structures of dominance. Specifically, we investigate how sentiment is assigned to the ingroup of collaborators (vs. the outgroup of scholars one has never co-authored with) and to the dominant group of scholars with high h-indices (vs. the less dominant group of scholars with lower h-indices).
@@ II.1 Citation Sentiment is Most Favorable to Collaborators
To examine how sentiment might reflect ingroup/outgroup relations, we measured the collaboration distance: the geodesic distance in the collaboration network between the citer and the citee [43] at the year of citation. We observed that people are most favorable to themselves and collaborators, and most critical to scholars with whom they have not collaborated (Fig. 2A). In subsequent analyses, we removed self-citations to focus on community practices. Note that collaboration distance likely tracks both social distance and intellectual distance. What happens when intellectual distance is held fixed, and social distance changes? To take an initial step toward addressing this question, we isolate citer-citee pairs who would eventually become collaborators. Just prior to collaborating, we find that sentiment is overwhelmingly critical, whereas just after collaborating, critical sentiment plummets and favorable sentiment rises (Fig. 2B). This pattern suggests that the social act of collaboration may drive a transformation from critical to favorable sentiment among scholars. We also notice that citer-citee pairs that will collaborate within the next four years tend to display significantly greater critical sentiment ($47.1\%$ more than expected) than citer-citee pairs that will not collaborate ($1.2\%$ more than expected; two-sided Welch’s $t$-test: $t=88.4$, $p<0.0001$, $df=1003.9$). The presence of this dramatic increase could indicate that critical language in citation may be an early predictor of future collaboration.
Figure 4: Critical citation sentiment varies across disciplines. (A) Mean critical sentiment in a discipline toward collaborators versus non-collaborators. (B) Mean critical sentiment as a function of the benchwork score of the citing paper’s discipline. We define the benchwork score as the estimated proportion of wet lab papers (as opposed to dry lab papers) in a discipline. Orange corresponds to the sentiment toward non-collaborators, cyan to the sentiment toward collaborators, and grey to the sentiment toward both. Lines are fit using weighted least squares regression; shaded regions represent $95\%$ confidence intervals; $s$ reports the slope of the regression. (C) Mean critical sentiment as a function of the synthesis score of a citing paper’s discipline. We define synthesis as the proportion of review papers in a discipline. (D) Relationship between the mean critical sentiment and bias. We define bias as the difference between sentiments toward non-collaborators and collaborators. (E) Bias in critical sentiment as a function of a discipline’s benchwork score. (F) Bias in critical sentiment as a function of disciplinary synthesis.
@@ II.3 Citation Sentiment Tracks Disciplinary Differences & Practices
Figure 5: Favorable citation sentiment varies across disciplines. (A) Mean favorable sentiment in a discipline toward collaborators versus non-collaborators. (B) Mean favorable sentiment as a function of the benchwork score of the citing paper’s discipline. We define benchwork score as the estimated proportion of wet lab papers (as opposed to dry lab papers) in a discipline. Orange corresponds to the sentiment toward non-collaborators, cyan to the sentiment toward collaborators, and grey to the sentiment toward both. Lines are fit using weighted least squares regression; shaded regions represent $95\%$ confidence intervals; $s$ reports the slope of the regression. (C) Mean favorable sentiment as a function of the synthesis score of a citing paper’s discipline. We define synthesis as the proportion of review papers in a discipline. (D) Relationship between the mean favorable sentiment and bias. We define bias as the difference between sentiments toward non-collaborators and collaborators. (E) Bias in favorable sentiment as a function of a discipline’s benchwork score. (F) Bias in favorable sentiment as a function of disciplinary synthesis.
The previous two sections suggest that citation sentiment tracks social factors at the level of inter-personal differences (h-index) and relationships (collaboration). We expand out from dyads to larger social groups in the form of scientific disciplines, with their own broader social norms, cultures [44, 45], and identities [46]. Moreover, as so-called communication complexes [47] with their own rhetoric [48], writing styles [49], and communicative processes [47], we expect disciplines to differ in citation sentiment and in the prevalence of critical vs. favorable sentiment to collaborators and strangers. Consistent with our expectations, we see that critical sentiment varies by discipline (Fig. 4A), with most disciplines displaying less critical sentiment to collaborators than to non-collaborators (Fig. 4A, the sector above the diagonal unity line). Interestingly, we observe that the degree of critical sentiment in a discipline overall is positively correlated with sentiment bias; that is, the more critical the discipline in general, the more likely the discipline will more critically cite non-collaborators than collaborators (Fig. 4D; linear regression slope $s=0.41$, $95\%$ CI $(0.17,0.65)$, adjusted $R^{2}=0.300$; one-sided F-test $F(1,25)=12.15$, $p<0.01$; $n=27$; all disciplinary results except for brilliance result share the same sample size). Similarly, we see that favorable sentiment varies by discipline (Fig. 5A), with all disciplines displaying more favorable sentiment to collaborators than to non-collaborators (Fig. 5A, the sector below the diagonal unity line). The degree of favorable sentiment in a discipline overall is significantly correlated with sentiment bias (Fig. 5D; slope $s=0.61$, $95\%$ CI $(0.37,0.84)$, adjusted $R^{2}=0.511$, $F(1,25)=28.19$, $p<0.0001$).
@@
We expect disciplines offering lower-level explanations from benchwork to have less critical sentiment because that scale of explanation is more accepted and valued. By contrast, we expect disciplines offering higher-level explanations to more frequently resort to the use of sentiment to signal valuation of the scientific work. Consistent with our expectations, we observe a significant negative trend between critical sentiment and the benchwork score (linear regression slope $s=-92.20$, $95\%$ CI $(-136.70,-47.70)$, adjusted $R^{2}=0.398$, $F(1,25)=18.21$, $p<0.001$; Fig. 4B, grey). Interestingly, the effect holds both when citing non-collaborators ($s=-97.07$, $95\%$ CI $(-145.39,-48.75)$, $R^{2}=0.383$, $F(1,25)=17.12$, $p<0.001$) and when citing collaborators ($s=-61.78$, $95\%$ CI $(-97.73,-25.84)$, $R^{2}=0.307$, $F(1,25)=12.53$, $p<0.01$). Similarly, we observe a significant negative trend between favorable sentiment and the benchwork score (linear regression slope $s=-32.71$, $95\%$ CI $(-48.72,-16.69)$, adjusted $R^{2}=0.391$; one-sided F-test $F(1,25)=17.69$, $p<0.001$; Fig. 5B, grey). Again the effect holds when citing non-collaborators ($s=-34.7$, $95\%$ CI $(-50.90,-18.50)$, $R^{2}=0.415$, $F(1,25)=19.46$, $p<0.001$), but is significantly mitigated when citing collaborators ($R^{2}=-0.001$, $F(1,25)=0.97$, $p=0.33$). Further, we observe a significantly negative trend between the benchwork score and bias in favorable sentiment towards collaborators (slope $s=-27.87$, $95\%$ CI $(-42.00,-13.74)$, adjusted $R^{2}=0.374$, $F(1,25)=16.51$, $p<0.001$; Fig. 5E). This negative trend is not seen for critical sentiment (linear fit and one-sided F-test between benchwork score and critical sentiment bias $R^{2}=0.058$, $F(1,25)=2.60$, $p=0.12$; Fig. 4E).
@@
We define synthesis to be the proportion of review papers in a discipline. Consistent with our expectations, we observe a significant negative correlation between critical sentiment and synthesis (linear regression slope $s=-307.97$, $95\%$ CI $(-407.70,-208.23)$, adjusted $R^{2}=0.603$; one-sided F-test $F(1,25)=40.45$, $p<0.0001$; Fig. 4C, grey). The effect holds when citing non-collaborators ($s=-330.24$, $95\%$ CI $(-434.39,-226.10)$, $R^{2}=0.616$, $F(1,25)=42.65$, $p<0.0001$), but is somewhat mitigated when citing collaborators ($s=-181.40$, $95\%$ CI $(-285.11,-77.70)$, $R^{2}=0.315$, $F(1,25)=12.98$, $p<0.01$). Further, we observe that disciplines with greater synthesis display less bias in critical sentiment towards non-collaborators, and disciplines with less synthesis display more bias (linear fit and one-sided F-test between synthesis and critical sentiment bias $s=-127.72$, $95\%$ CI $(-224.11,-31.33)$, $R^{2}=0.199$, $F(1,25)=7.45$, $p<0.05$; Fig. 4F). Similarly, we observe a significant negative correlation between favorable sentiment and synthesis (linear regression slope $s=-114.84$, $95\%$ CI $(-149.75,-79.92)$, adjusted $R^{2}=0.633$; one-sided F-test $F(1,25)=45.88$, $p<0.0001$; Fig. 5C, grey). Again the effect holds when citing non-collaborators ($s=-116.28$, $95\%$ CI $(-153.00,-79.56)$, $R^{2}=0.615$, $F(1,25)=42.54$, $p<0.0001$), and is mitigated when citing collaborators ($R^{2}=0.100$, $F(1,25)=3.88$, $p=0.06$). We observe a significant negative relationship between synthesis and bias in favorable sentiment towards collaborators (slope $s=-90.24$, $95\%$ CI $(-126.48,-54.01)$, adjusted $R^{2}=0.493$, $F(1,25)=26.31$, $p<0.0001$; Fig. 5F). Note that our null model (see Methods) accounts for the fact that review papers in general express less sentiment ($35.64\%$ favorable, $6.79\%$ critical) compared to research papers ($44.71\%$ favorable, $8.37\%$ critical). This result also holds before the sentiment aggregation step in our methods (see Supplement).